A standing wave is set up in string 120 cm long. The string vibrates in four segments under a frequency of 120 Hz. Calculate the wavelength and the fundamental frequency of the string.


1
Expert's answer
2021-05-10T02:42:13-0400

Length of string(l)=120cm=1.20m

String vibrate in four segment. Mean numbers of loop is n=4

Wavelength

"\lambda n=2l/n"

=2*1.20/4=0.60m

Fundamental frequency=120/4=30Hz